I need help with two more problems. They do not come from a book. I know the answers but am lost on the way to do them. I have a big test tomorrow and need to understand these a little better. Will you please show me the steps? Thank you very much.

1.) Solve for x. x^2-x+4=0
Answer is ½+sqrt15/2i, ½-sqrt15/2i

2.) Solve (x^2-6x)^2 - 11(x^2-6x)-80=0
I think it may be let u=x^2-6x, but was not sure. The answer is 5, 1, 8, -2

Answer by scott8148(6628)

1) this uses the quadratic formula
__ for ax^2+bx+c=0 __ x=(-b plus or minus sqrt(b^2-4ac))/2a

2) u^2-11u-80=0
__ factoring __ (u-16)(u+5)=0 __ u=16 and u=-5

substituting __ x^2-6x=16 __ x^2-6x-16=0
__ factoring __ (x-8)(x+2)=0 __ x=8 and x=-2

substituting __ x^2-6x=-5 __ x^2-6x+5=0
__ factoring __ (x-5)(x-1)=0 __ x=5 and x=1